How do I calculate?At the beginning of the year, Infodeo established its predetermined overhead rate for movies produced during the year by usinReg 1 Req 2 and 3 Reg 4 Enter the overhead costs incurred and the amounts applied during the year using the predetermined ove
Answer #1

SOLUTION

1. Predetermined overhead rate = Estimated overhead / Estimated direct labor cost * 100

= $1,833,000 / 470,000 = 390%

2&3. Applied overhead costs = Direct labor cost * Predetermined overhead rate

= 244,000 * 390% = $951,600

Factory overhead

Actual overhead 940,200 Applied overhead 951,600
Factory overhead (Over applied)) 11,400

Over applied overhead = Applied overhead - Actual overhead

= 951,600 - 940,200 = 11,400

4. Journal entry-

Factory overhead 11,400
Cost of goods sold 11,400
(To assign over applied overhead to cost of goods sold)